# Multi-View Camera System for Variant-Aware Autonomous Vehicle Inspection and Defect Detection

### v1

@@ Abstract
I Introduction
@@ II Related Work
II-A Multi-camera perception in autonomous vehicles
@@
II-B Vision‑based quality inspection and control
@@
II-C Information fusion and defect detection
Classical fusion in automotive inspection often combines LiDAR and cameras at the raw/data level [25, 26], which is powerful yet sensitive to calibration and noise. Decision‑level fusion, aggregating per‑sensor inferences, offers lighter, noise‑tolerant deployment [27]. We follow this latter paradigm: view‑aware detectors produce standardized evidence that a semantic rule engine evaluates against a VIN‑conditioned manifest. For defects, real vehicles exhibit long‑tail, rare patterns where large labeled datasets are scarce [10, 8]. Detectors are therefore appropriate; in our pipeline, a segmentation module provides data-efficient damage cues that seamlessly integrate with the decision layer.
III Proposed Method: The AVI System
TABLE I: AVI Camera Layout and Key Inspection Targets.
@@ III Proposed Method: The AVI System
III-A System Architecture and Camera Layout
@@
The camera configuration and its primary inspection targets are detailed in Table I.
@@
III-B View-Aware Fusion and Semantic Decision Engine
@@ IV-A Dataset
IV-B Tasks & Evaluation Metrics
@@ IV-D Quantitative Results
IV-D1 Stage 1: Vehicle Identification and Branding Detection
We detect logo, mascot, and front_grille as inputs to OCR and ICE/EV classification. Table II reports Precision(P), Recall(R), Mean Average Precision(mAP)@50, and F1.
TABLE II: Stage 1: object detection for branding cues.
@@ IV-D1 Stage 1: Vehicle Identification and Branding Detection
IV-D2 Stage 2: Variant‑Specific Feature Detection
We detect roof rails, rear antenna, and rear wiper from top/rear views and side cues for wheel type. Results in Table III again favor YOLOv8.
TABLE III: Stage 2: variant feature detection.
@@ IV-D2 Stage 2: Variant‑Specific Feature Detection
IV-D3 Stage 3: Surface Anomaly Segmentation
We compare instance‑segmentation backbones for scratches/dents (Table IV); YOLOv8‑Seg attains the best mAP and mIoU. The BB stands for bounding box, in (Table IV).
TABLE IV: Stage 3: scratch/dent segmentation.
@@ IV-D3 Stage 3: Surface Anomaly Segmentation
IV-E Ablation Study
We quantify the impact of multi‑view design and the damage module (Table V). Single‑view baselines lack feature coverage; removing segmentation reduces defect recall despite lower latency. Here, Feat. Cov. (%) denotes Feature Coverage, Def. Det. (%) represents Defect Detection, Lat. (ms) refers to Latency, and Verif. Acc. (%) corresponds to Verification Accuracy.
TABLE V: Ablations: multi‑view value and segmentation contribution.
@@ IV-E Ablation Study
The multi‑view rig is necessary for complete coverage; the segmentation branch raises defect detection by $\approx$14–18 pp with an acceptable latency trade‑off, yielding a balanced system suitable for deployment.
V Conclusion

### v2

# Multi-View Camera Perception System for Variant-Aware Autonomous Vehicle Inspection and Defect Detection

@@ Abstract
keywords:
Semantic occupancy prediction, Multi-sensor data fusion, Vehicle Inspection, Defect Detection, Quality Control, Scene understanding, Autonomous driving, Feature fusion
1 Introduction
@@ 2 Related Work
2.1 Multi-camera perception in autonomous vehicles
@@
2.2 Vision‑based quality inspection and control
@@
2.3 Information fusion and defect detection
Classical fusion in automotive inspection often combines LiDAR and cameras at the raw/data level [25, 26], which is powerful yet sensitive to calibration and noise. Decision‑level fusion, aggregating per‑sensor inferences, offers lighter, noise‑tolerant deployment [27, 28]. We follow this latter paradigm: view‑aware detectors produce standardized evidence that a semantic rule engine evaluates against a VIN‑conditioned manifest. For defects, real vehicles exhibit long‑tail, rare patterns where large labeled datasets are scarce [10, 8]. Detectors are therefore appropriate; in our pipeline, a segmentation module provides data-efficient damage cues that seamlessly integrate with the decision layer.
3 Proposed Method: The AVI System
Table 1: AVI Camera Layout and Key Inspection Targets.
@@ 3 Proposed Method: The AVI System
3.1 System Architecture and Camera Layout
@@
The camera configuration and its primary inspection targets are detailed in Table 1.
@@
3.2 View-Aware Fusion and Semantic Decision Engine
@@ 4.1 Dataset
4.2 Tasks & Evaluation Metrics
@@ 4.4 Quantitative Results
4.4.1 Stage 1: Vehicle Identification and Branding Detection
We detect logo, mascot, and front_grille as inputs to OCR and ICE/EV classification. Table 2 reports Precision(P), Recall(R), Mean Average Precision(mAP)@50, and F1.
Table 2: Stage 1: object detection for branding cues.
@@ 4.4.1 Stage 1: Vehicle Identification and Branding Detection
4.4.2 Stage 2: Variant‑Specific Feature Detection
We detect roof rails, rear antenna, and rear wiper from top/rear views and side cues for wheel type. Results in Table 3 again favor YOLOv8.
Table 3: Stage 2: variant feature detection.
@@ 4.4.2 Stage 2: Variant‑Specific Feature Detection
4.4.3 Stage 3: Surface Anomaly Segmentation
We compare instance‑segmentation backbones for scratches/dents (Table 4); YOLOv8‑Seg attains the best mAP and mIoU. The BB stands for bounding box, in (Table 4).
Table 4: Stage 3: scratch/dent segmentation.
@@ 4.4.3 Stage 3: Surface Anomaly Segmentation
4.5 Ablation Study
We quantify the impact of multi‑view design and the damage module (Table 5). Single‑view baselines lack feature coverage; removing segmentation reduces defect recall despite lower latency. Here, Feat. Cov. (%) denotes Feature Coverage, Def. Det. (%) represents Defect Detection, Lat. (ms) refers to Latency, and Verif. Acc. (%) corresponds to Verification Accuracy. The multi‑view rig is necessary for complete coverage; the segmentation branch raises defect detection by $\approx$14–18 pp with an acceptable latency trade‑off, yielding a balanced system suitable for deployment.
Table 5: Ablations: multi‑view value and segmentation contribution.
@@ 4.5 Ablation Study
5 Discussion
5.1 Bridging the Perception-Reasoning Gap in Industrial Autonomy
While recent advances in deep learning have solved the what is it problem (detection), industrial autonomy requires answering the is it correct problem (verification). Standard object detectors lack the contextual awareness to distinguish between a valid missing part (e.g., a base-model vehicle lacking a sunroof) and a defect (e.g., a premium model missing a sunroof). Our work demonstrates that relying solely on data-driven perception is insufficient for variant-aware inspection. The AVI system bridges this gap through its Semantic Decision Engine. By decoupling visual detection from logic to validating standardized evidence against a VIN-derived configuration, we transform probabilistic detections into deterministic pass/fail decisions. This architecture suggests a blueprint for future autonomous quality control systems: perception modules should act as sensors that feed a deterministic reasoning layer, rather than attempting to learn complex business logic end-to-end.
5.2 The Necessity of Multi-View Fusion for Robustness
A critical critique of autonomous visual inspection is the reliance on single-viewpoint sensors, which fail under occlusion. Our ablation study (Table 5) empirically quantifies this feature coverage risk. Single views (Front T1 or Rear T3) achieve only 28-71% feature coverage, leading to unacceptable verification accuracy (88-93%). The No Segmentation baseline highlights another trade-off: while excluding the segmentation module reduces latency by $\approx$70 ms, it causes a catastrophic drop in defect detection (from 86.5% to $\approx$70%)5. This confirms that in autonomous industrial settings, sensor redundancy (11 cameras) and task specialization (segmentation vs. detection) are not inefficiencies, but rigid requirements for reliability.
5.3 Deployability vs. Complexity in Industry 4.0
Unlike academic benchmarks that prioritize mAP improvements, real-world deployment is constrained by the tact time of production lines. The AVI system sustains a throughput of 3.3 vehicles/minute ($\approx$18 seconds/vehicle) with a processing latency of 285 ms. This performance is achieved by avoiding heavy, monolithic transformers in favor of optimized, task-specific backbones (YOLOv8, EfficientNet). While newer architectures may offer marginal accuracy gains, our results show that the primary bottleneck in industrial autonomy is not detector precision, but the integration of multi-modal evidence (OCR, classification, segmentation) into a unified decision. The choice of established backbones ensures stability and ease of retraining, a critical factor for factory adoption where data for new defects is often scarce.
5.4 Limitations and Future Directions
While our View-Aware Fusion layer effectively standardizes evidence, it currently relies on a static camera-to-task mapping. Future iterations could explore adaptive view attention, where the system dynamically queries different cameras based on confidence scores or detected ambiguities (e.g., zooming in on a scratch if the L1 view is occluded). Additionally, while the VIN-conditioned logic handles known variants, extending the system to open-world anomaly detection, identifying completely unknown foreign objects without prior training, remains a frontier for autonomous inspection research.
6 Conclusion
@@
[…]
Acknowledgements
7 Declarations
  • Funding
  • Author contribution
  • Conflict of interest
  • Data availability